# Towards AI-Safety-by-Design: A Taxonomy of Runtime Guardrails in Foundation Model based Systems

### v1

@@ VI Conclusion and Future Work
To better understand runtime guardrails in FM-based systems, this paper developed and presented a comprehensive guardrail taxonomy and provided a comparative study of its components based on a systematic literature review. Our proposed taxonomy categorizes guardrails based on their fundamental objectives, essential quality attributes, performed actions, targeted aspects, covered scopes, employed rules, autonomy, modalities, and underlying techniques.
Our key findings emphasize that the effective integration of guardrails mitigates risk such as biases, harmful content, and unintended behaviors in FM-based systems. Practitioners can use the proposed taxonomy to improve the design of guardrails to ensure the behavior and decisions of FM-based systems are responsible and safe. Moreover, the taxonomy provides a theoretical framework that researchers and policymakers can use to develop guidelines and standards that promote AI-safety-by-design in FM-based systems. In the future, we plan to develop a guardrails service for FM-based agents, integrating various design options outlined in this taxonomy.

### v2

@@ VI Conclusion and Future Work
To better understand runtime guardrails in FM-based agents, this paper developed and presented a comprehensive guardrail taxonomy and provided a comparative study of its components based on a systematic literature review guided by the Swiss Cheese Model for AI safety. Our proposed taxonomy categorizes guardrails based on their fundamental objectives, essential quality attributes, performed actions, targeted aspects, covered scopes, employed rules, autonomy, modalities, and underlying techniques.
Our key findings emphasize that the effective integration of guardrails mitigates risk such as biases, harmful content, and unintended behaviors in FM-based agents. Practitioners can use the proposed taxonomy to improve the design of guardrails to ensure the behavior and decisions of FM-based agents are responsible and safe. Moreover, the taxonomy provides a theoretical framework that researchers and policymakers can use to develop guidelines and standards that promote AI-safety-by-design in FM-based agents.
In the future, we plan to develop a multilayer architecture for guardrails in FM-based agents, integrating various design options outlined in this taxonomy. Additionally, we aim to evaluate the effectiveness of multi-layered guardrails in real-world agent deployments, further refining the taxonomy and contributing to best practices in AI safety.